OpenStudy (anonymous):

Isolate H: V=1/3*pi*r^2h Please show me the steps.. I am a bit lost and want to get this right.

OpenStudy (zepp):

\[\large V = \frac{1}{3}\pi r^2h\] Like this?

OpenStudy (anonymous):

yeahhh!!

OpenStudy (anonymous):

but i need to isolate H

OpenStudy (zepp):

First step, take out the \[\frac{1}{3}\] By dividing by 1/3 both side \[\large V = \frac{1}{3}\pi r^2 h\]\[\large V \div \frac{1}{3} =\pi r^2 h\]

OpenStudy (zepp):

\[V \div \frac{1}{3}= 3V\]

OpenStudy (anonymous):

doesn't it go to 3/1 instead?? which is the same as 3...

OpenStudy (zepp):

\[\large 3V = \pi r^2 h\] Move \[\pi r^2\] \[\large 3V \div \pi r ^2 =h\] \[\large \frac{3V}{\pi r ^2} =h\]

OpenStudy (anonymous):

thats what i got. okay thanks:)

OpenStudy (zepp):

Np ;)